# Cell as Point: One-Stage Framework for Efficient Cell Tracking

### v2

@@ B.2 DeepCell Dataset
Table 5: Hyperparameter Configuration for CAP.
Metrics HeLa PC-3 3T3 RAW264
KIT-GE CAP KIT-GE CAP KIT-GE CAP KIT-GE CAP
$\mathrm{ES}$downarrow 6 3 29 1 128 107 37 67
$\mathrm{EA}$downarrow 57 82 56 11 783 949 585 1098
$\mathrm{ED}$downarrow 1 1 0 0 14 2 106 12
$\mathrm{ESM}$downarrow 17 0 0 0 10 4 24 17
$\mathrm{FP}$downarrow 70 60 13 35 256 250 718 647
$\mathrm{FN}$downarrow 67 70 1 7 698 709 603 898
$\mathrm{TRA}$uparrow 0.923 0.926 0.937 0.952 0.820 0.832 0.876 0.864
$\mathrm{IT}$downarrow 14.4 1.3 9.1 1.1 19.1 2.2 96.7 2.8
Table 6: Comparison Experiment with Baseline. Here is a complementary for Tab. 1 in Sec. 4.2.
Appendix C Experiment Details
C.1 Training and Optimizer
By utilizing the pre-trained CoTracker [40] checkpoint, we freeze the CNN feature extractor $\phi$ and use the AdamW optimizer to train the proposed CAP model. We present hyperparameter configuration for the optimizer in Tab. 5.
C.2 Evaluation Metrics
We follow the evaluation metrics of the operations to correct the situations of merged cells, missing links, wrong links, links with wrong semantics, false positives, and false negatives: $\mathrm{EdgeSplit}$ ($\mathrm{ES}$), $\mathrm{EdgeAdd}$ ($\mathrm{EA}$), $\mathrm{EdgeDelete}$ ($\mathrm{ED}$), $\mathrm{EdgeSemantic}$ ($\mathrm{ESM}$), $\mathrm{FP}$, and $\mathrm{FN}$. $\mathrm{TRA}$ measure [52, 74] is calculated from AOGM:
$\mathrm{TRA}=1-\frac{\mathrm{min}(AOGM,AOGM_{0})}{AOGM_{0}},$ (9)
@@ C.2 Evaluation Metrics
$w_{\mathrm{ESM}}\mathrm{ESM}+w_{\mathrm{FP}}\mathrm{FP}+w_{\mathrm{FN}}\mathrm{FN}.$ (10)
@@
AOGM is the weighted sum of the executed operations considered as the cost of transforming ($\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, $\mathrm{ESM}$, $\mathrm{FP}$, and $\mathrm{FN}$) the computed graph into the reference one. and $\mathrm{TRA}$ used in the ISBI cell tracking challenge. The details of $\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, $\mathrm{ESM}$, $\mathrm{FP}$, and $\mathrm{FN}$ are elaborated in the following:
AEG Sampling HeLa PC-3 3T3 RAW264
@@
0.898 0.908 0.883 0.921 0.878 0.952 0.924 0.829 0.804 0.828 0.843
0.962 0.958 0.921 0.900 0.891 0.980 0.924 0.832 0.824 0.868 0.900
Table 7: Effectiveness of AEG Sampling. AEG sampling is built to effectively balance the cell event (division) in a training sequence.
Attention HeLa PC-3 3T3 RAW264
@@
Stride HeLa PC-3 3T3 RAW264
@@
$\mathrm{\mathbf{ES}}$ is the operation for non-split vertices that the computed markers with more than one reference marker assigned.
$\mathrm{\mathbf{EA}}$ is the operation for missing links that the cell lineage relations are not indicated.
$\mathrm{\mathbf{ED}}$ is the operation for redundant links that the fake cell lineage relations are indicated.
$\mathrm{\mathbf{ESM}}$ is the operation for altering the edge semantics, that is, the difference in semantics between the reference graph and the induced subgraph.
$\mathrm{\mathbf{FP}}$ is the situation in the extra detected objects (i.e., computed markers without any reference marker assigned).
$\mathrm{\mathbf{FN}}$ is the situation in the missed objects (i.e., reference markers not assigned to any computed marker).
Appendix D Additional Baseline Results
Tab. 6 provides additional baseline results as a complementary analysis to Tab. 1 in Sec. 4.2., further demonstrating the effectiveness of our CAP framework. It shows the performance comparison of our proposed CAP framework against KIT-GE across four datasets, HeLa, PC-3, 3T3, and RAW264, in the DeepCell. CAP consistently achieves higher TRA (tracking accuracy) scores than KIT-GE in HeLa, PC-3, and 3T3 datasets, demonstrating its superior tracking capability. Furthermore, CAP significantly reduces IT (inference time) across all datasets, being 10×10$\times$10 × to $\times8$ faster than KIT-GE, highlighting its efficiency. Overall, CAP delivers superior tracking accuracy with significantly lower inference time, reinforcing its effectiveness as a one-stage tracking solution compared to the multi-stage KIT-GE framework.
Appendix E Additional Ablations
Following the initial assessment of adaptive event-guided (AEG) sampling, cell joint tracking structure, and feature stride on the ISBI dataset in Sec. 4.3, we extended our ablation studies to the DeepCell dataset to validate the generalizability and effectiveness of these components in our CAP framework. The findings, presented in Tab. 7Tab. 8, and Tab. 9, underscore the significant impact of each component on the framework’s performance across various cell types and conditions.
E.1 Effectiveness of AEG Sampling
Tab. 7 shows that adaptive AEG sampling effectively balances cell event occurrences during training sequences, which enables AEG sampling to generally improve performance across all cell types. For instance, in the HeLa dataset, AEG sampling boosts the scores from $0.898$ to $0.962$ in the best case, indicating a substantial enhancement in the model’s ability to handle imbalanced data effectively. This improvement is consistent across other cell types like PC-3 and RAW264, affirming AEG sampling’s role in optimizing event-centric learning.
E.2 Effectiveness of Cell Joint Tracking
As represented in the main manuscript, we replace the cross-trajectory attention layers with only time attention layers to keep the model size. As illustrated in Tab. 8, the integration of both time and joint attention mechanisms in cell joint tracking significantly enhances tracking accuracy. The dual attention strategy not only improves temporal coherence but also reinforces the relational context between trajectories, evident from the increased scores in datasets such as HeLa, where the combined attention boosts the performance from $0.853$ to $0.962$ in the best cases. This enhancement highlights the importance of sophisticated tracking mechanisms in complex cellular environments.
E.3 Ablation on Feature Stride
The impact of feature stride on model accuracy is depicted in Tab. 9. Reducing the stride from $8$ to $4$ leads to a marked improvement in the resolution of feature maps, which in turn enhances the model’s predictive accuracy. For example, in the HeLa dataset, reducing the stride results in an increase from $0.884$ to $0.962$ in the best scenarios. This finding suggests that higher resolution features are critical for accurate cell tracking, particularly in datasets with intricate cell dynamics and small cell sizes like 3T3 and RAW264.

### v3

@@ 4.1.1 Datasets
ISBI CTC.
Fig. 5 illustrates datasets we leveraged, U373, GOWT1, and Huh7, with the comparison between tracking (TRA) GT mask, segmentation (SEG) GT, and segmentation (SEG) ST if available. We observe that SEG GT and ST are much more refined and high-quality than TRA GT masks, requiring more resources and manpower. In terms of data annotation requirements, unlike previous works that rely on all available TRA GT masks, SEG GT, and ST (when available), CAP requires only the TRA GT masks, thereby substantially reducing the annotation demand. The detailed information on data collection and the characteristics of each dataset is described below:
  • Fluo-C2DL-Huh7 is a dataset of human hepatocarcinoma-derived cells expressing the fusion protein YFP-TIA-1. Cell events are captured by the Nikon Eclipse Ti2 microscope with a lens CFI Plan Apo Lambda 20x/0.75 in a duration of 15 minutes with $30$ frames for each sequence. Cells in this dataset exhibit a relatively low frequency of cell events; However, due to the lower frame rate, cells exhibit significantly larger movements within a sequence. Cells are relatively small and have high density in frames.
Implementation Details.
For the training process, we utilize $85$ training sequences from the DeepCell dataset and sequences $02$ from the ISBI dataset, using only tracking (TRA) ground truth (GT). During each iteration, we crop sequences of $T_{s}=24$ frames with a sliding window length of $l_{\text{slide}}=8$. We use pre-trained checkpoints from CoTracker [43], which was trained solely on synthetic RGB data. Since our input images are grayscale, we replicate the single channel three times to match the required RGB format. We use an end-to-end trained $2$-dimensional CNN architecture as the encoder to extract tracking features $F$. Specifically, it integrates bilinear interpolation and utilizes it to extract $S=4$ different scaled image features with $128$-channel output features concatenated to optimized dynamic feature extraction. While training the transformer, we follow the pre-trained checkpoint that downsamples the resolution of frames to $384\times 512$ and choose $S=4$, $\Delta=3$ to calculate 4D multi-scaled correlation volumes. As for the optimization during the training, we present an optimizer configuration in Table 1.
@@
Table 1: Optimizer Configuration for CAP.
As for the testing (evaluation) process, we preprocess the first frame for each sequence using [70] to select query points $Q$ and start the process of window rolling. The evaluation sequence $S_{\text{infer}}$ follows the native resolution ($384\times 512$), which the dataset is trained on. We set $l_{\text{win}}=100$, which is mentioned in Sec. 3.2.3. We follow the $\mathrm{TRA}$ evaluation metrics [58, 41] of the operations provided by ISBI [57, 80] and CTMC [4] to correct the situations of merged cells, missing links, wrong links, links with wrong semantics, false positives, and false negatives: $\mathrm{EdgeSplit}$ ($\mathrm{ES}$), $\mathrm{EdgeAdd}$ ($\mathrm{EA}$), $\mathrm{EdgeDelete}$ ($\mathrm{ED}$), $\mathrm{EdgeSemantic}$ ($\mathrm{ESM}$), $\mathrm{FP}$, and $\mathrm{FN}$. $\mathrm{TRA}$ lies in the range $[0,1]$ where a higher score corresponds with a better performance, and is calculated by:
$\mathrm{TRA}=$ $1-\frac{\mathrm{min}(AOGM,AOGM_{0})}{AOGM_{0}},\text{ where}$ (7)
@@
$w_{\mathrm{ESM}}\mathrm{ESM}+w_{\mathrm{FP}}\mathrm{FP}+w_{\mathrm{FN}}\mathrm{FN}.$
@@
Here $AOGM$ is the weighted sum of the executed operations considered as the cost of transforming ($\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, $\mathrm{ESM}$, $\mathrm{FP}$, and $\mathrm{FN}$) the computed graph into the reference one, and $AOGM_{0}$ is the $AOGM$ value for empty tracking results. The details of $\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, $\mathrm{ESM}$, $\mathrm{FP}$, and $\mathrm{FN}$ are elaborated as follows [^1] https://celltrackingchallenge.net/evaluation-methodology/:
Method Type HeLa
$\mathrm{EdgeSplit}$ $\downarrow$ $\mathrm{EdgeAdd}$ $\downarrow$ $\mathrm{EdgeDelete}$ $\downarrow$ $\mathrm{EdgeSemantic}$ $\downarrow$ $\mathrm{FP}$ $\downarrow$ $\mathrm{FN}$ $\downarrow$ $\mathrm{TRA}$ $\uparrow$ $\mathrm{IT}$ $\downarrow$
GNN [8] 22’ SEG+TRA 2 385 1 13 62 342 0.776 23.3
EmbedTrack [54] 22’ SEG+TRA 0 211 1 1 31 218 0.849 41.5
KIT-GE [70] 20’ SEG+TRA 6 57 1 17 70 67 0.923 14.4
Trackastra [26] 24’ SEG+TRA 1 240 1 1 124 185 0.825 11.8
Cell-TRACTR [64] 25’ SEG+TRA 6 209 1 1 715 144 0.816 30.3
CAP (Ours) TRA 3 82 1 0 60 70 0.926 1.3
Method Type PC-3
$\mathrm{EdgeSplit}$ $\downarrow$ $\mathrm{EdgeAdd}$ $\downarrow$ $\mathrm{EdgeDelete}$ $\downarrow$ $\mathrm{EdgeSemantic}$ $\downarrow$ $\mathrm{FP}$ $\downarrow$ $\mathrm{FN}$ $\downarrow$ $\mathrm{TRA}$ $\uparrow$ $\mathrm{IT}$ $\downarrow$
GNN [8] 22’ SEG+TRA 2 52 3 12 33 0 0.868 18.3
EmbedTrack [54] 22’ SEG+TRA 0 43 3 1 0 42 0.912 48.7
KIT-GE [70] 20’ SEG+TRA 29 56 0 0 13 1 0.937 9.1
Trackastra [26] 24’ SEG+TRA 2 233 2 1 155 190 0.317 8.9
Cell-TRACTR [64] 25’ SEG+TRA 1 26 1 0 146 8 0.913 33.7
CAP (Ours) TRA 1 11 0 0 35 7 0.952 1.1
Method Type 3T3
$\mathrm{EdgeSplit}$ $\downarrow$ $\mathrm{EdgeAdd}$ $\downarrow$ $\mathrm{EdgeDelete}$ $\downarrow$ $\mathrm{EdgeSemantic}$ $\downarrow$ $\mathrm{FP}$ $\downarrow$ $\mathrm{FN}$ $\downarrow$ $\mathrm{TRA}$ $\uparrow$ $\mathrm{IT}$ $\downarrow$
GNN [8] 22’ SEG+TRA 6 613 16 3 100 9 0.857 21.7
EmbedTrack [54] 22’ SEG+TRA 2 603 14 4 201 589 0.853 74.6
KIT-GE [70] 20’ SEG+TRA 128 783 14 10 256 698 0.820 19.1
Trackastra [26] 24’ SEG+TRA 10 2597 16 21 1511 1936 0.572 12.0
Cell-TRACTR [64] 25’ SEG+TRA 6 757 3 10 11834 213 0.727 65.0
CAP (Ours) TRA 53 849 2 4 189 509 0.854 2.2
Method Type RAW264
$\mathrm{EdgeSplit}$ $\downarrow$ $\mathrm{EdgeAdd}$ $\downarrow$ $\mathrm{EdgeDelete}$ $\downarrow$ $\mathrm{EdgeSemantic}$ $\downarrow$ $\mathrm{FP}$ $\downarrow$ $\mathrm{FN}$ $\downarrow$ $\mathrm{TRA}$ $\uparrow$ $\mathrm{IT}$ $\downarrow$
GNN [8] 22’ SEG+TRA 186 1451 89 21 239 1341 0.801 28.5
EmbedTrack [54] 22’ SEG+TRA 68 1475 74 28 669 1259 0.803 59.6
KIT-GE [70] 20’ SEG+TRA 37 585 106 24 718 603 0.876 96.7
Trackastra [26] 24’ SEG+TRA 14 4572 82 46 2810 3354 0.490 56.3
Cell-TRACTR [64] 25’ SEG+TRA 21 1346 11 12 16617 354 0.713 51.5
CAP (Ours) TRA 67 1098 12 17 647 898 0.864 2.8
Table 2: Comparison Experiments with Baselines. We compare CAP to existing cell tracking frameworks available on the HeLa, PC-3, 3T3, and RAW264 datasets in the DeepCell database.
$\mathrm{\mathbf{ES}}$ is the operation for non-split vertices that the computed markers with more than one reference marker assigned;
$\mathrm{\mathbf{EA}}$ is the operation for missing links that the cell lineage relations are not indicated;
$\mathrm{\mathbf{ED}}$ is the operation for redundant links that the fake cell lineage relations are indicated;
$\mathrm{\mathbf{ESM}}$ is the operation for altering the edge semantics, that is, the difference in semantics between the reference graph and the induced subgraph;
$\mathrm{\mathbf{FP}}$ is the situation in the extra detected objects;
$\mathrm{\mathbf{FN}}$ is the situation in the missed objects.
Baselines.
To evaluate the performance of CAP, we conducted a comparison experiment with baseline methods, cell-tracker-gnn (GNN) [8], EmbedTrack [54], KIT-GE [70], Trackastra [26], Cell-TRACTR [64], and EBB [45], which have publicly available code and utilize the evaluation system provided by ISBI [57, 80] and the acyclic oriented graphs matching (AOGM) measure [58]. For the DeepCell data, we follow the experiments [59] and the default training/testing subsets to report the result. For the ISBI CTC dataset, to control for variables and ensure fairness, we train these methods on sequence $02$ and evaluate on $01$, following [70]. Since the EBB baseline method does not provide pretrained weights for detection (U373 and GOWT1), we report the performance metrics as presented in its original paper[45]. We report the measures of metrics involved in the cell tracking, $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, $\mathrm{FP}$, $\mathrm{FN}$, $\mathrm{TRA}$, and the average inference time ($\mathrm{IT}$) to evaluate the cell tracking performance and the inference efficiency.
Method Type U373
$\mathrm{EdgeSplit}$ $\downarrow$ $\mathrm{EdgeAdd}$ $\downarrow$ $\mathrm{EdgeDelete}$ $\downarrow$ $\mathrm{EdgeSemantic}$ $\downarrow$ $\mathrm{FP}$ $\downarrow$ $\mathrm{FN}$ $\downarrow$ $\mathrm{TRA}$ $\uparrow$ $\mathrm{IT}$ $\downarrow$
GNN [8] 22’ SEG+TRA 214 371 0 0 135 25 0.771 19.7
EmbedTrack [54] 22’ SEG+TRA 0 44 0 7 211 32 0.931 123.6
KIT-GE [70] 20’ SEG+TRA 3 50 0 1 197 25 0.939 22.0
Trackastra [26] 24’ SEG+TRA 3 50 0 21 182 25 0.938 20.3
EBB [45] 25’ DET+TRA - - - - - - 0.967 -
CAP (Ours) TRA 0 0 0 0 119 0 0.985 2.6
Method Type GOWT1
$\mathrm{EdgeSplit}$ $\downarrow$ $\mathrm{EdgeAdd}$ $\downarrow$ $\mathrm{EdgeDelete}$ $\downarrow$ $\mathrm{EdgeSemantic}$ $\downarrow$ $\mathrm{FP}$ $\downarrow$ $\mathrm{FN}$ $\downarrow$ $\mathrm{TRA}$ $\uparrow$ $\mathrm{IT}$ $\downarrow$
GNN [8] 22’ SEG+TRA 49 140 1 13 5 28 0.958 67.2
EmbedTrack [54] 22’ SEG+TRA 39 119 0 0 21 40 0.961 171.6
KIT-GE [70] 20’ SEG+TRA 45 111 0 1 6 32 0.966 39.4
Trackastra [26] 24’ SEG+TRA 47 140 0 1 4 36 0.965 27.6
EBB [45] 25’ DET+TRA - - - - - - 0.991 -
CAP (Ours) TRA 18 120 2 2 75 46 0.960 7.1
Method Type Huh7
$\mathrm{EdgeSplit}$ $\downarrow$ $\mathrm{EdgeAdd}$ $\downarrow$ $\mathrm{EdgeDelete}$ $\downarrow$ $\mathrm{EdgeSemantic}$ $\downarrow$ $\mathrm{FP}$ $\downarrow$ $\mathrm{FN}$ $\downarrow$ $\mathrm{TRA}$ $\uparrow$ $\mathrm{IT}$ $\downarrow$
GNN [8] 22’ SEG+TRA 34 85 2 835 36 4 0.889 27.1
EmbedTrack [54] 22’ SEG+TRA 0 733 0 5 1392 636 0.187 125.0
KIT-GE [70] 20’ SEG+TRA 34 77 0 0 41 0 0.970 25.7
Trackastra [26] 24’ SEG+TRA 34 84 2 5 34 5 0.964 18.2
EBB [45] 25’ DET+TRA - - - - - - 0.626 -
CAP (Ours) TRA 31 71 0 1 101 1 0.966 3.2
Table 3: Comparison Experiments of ISBI CTC Dataset. We compare CAP to existing cell tracking frameworks available on the datasets PhC-C2DH-U373, Fluo-N2DH-GOWT1, and Fluo-C2DL-Huh7.
4.2 Results and Analysis
We performed comparison experiments of our method, CAP, against baseline approaches, including GNN, EmbedTrack, KIT-GE, Trackastra, Cell-TRACTR, and EBB across the DeepCell datasets (HeLa, PC-3, 3T3, and RAW264) and ISBI datasets (U373, GOWT1, and Huh7). Evaluation is based on both cell-tracking accuracy-related metrics ($\mathrm{TRA}$ derived from $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, $\mathrm{FP}$, and $\mathrm{FN}$) and inference efficiency ($\mathrm{IT}$).
Table 2 demonstrates the quantitative comparison results on DeepCell datasets. On HeLa and PC-3, CAP achieves the highest $\mathrm{TRA}$ scores of $0.926$ and $0.952$, surpassing prior works while requiring only $1.3$s and $1.1$s inference time, respectively. This highlights CAP’s ability to deliver state-of-the-art accuracy with minimal computational cost. On RAW264, CAP attains a competitive $\mathrm{TRA}$ of $0.864$ with just $2.8$s inference, demonstrating a $3-20\times$ reduction in runtime compared to other methods. Even on the challenging 3T3 dataset, CAP achieves a strong $\mathrm{TRA}$ of $0.854$, comparable to the best baseline ($0.857$), but with an inference time of only $2.2$s versus up to $65.0$s for competing methods. As for the ISBI CTC datasets, as shown in Table 3, on U373, CAP achieves the best overall performance, recording zero errors across tracking structural metrics and achieving a $\mathrm{TRA}$ of $0.985$, with the lowest inference time of $2.6$s. For GOWT1, CAP achieves a $\mathrm{TRA}$ of $0.960$, slightly below KIT-GE ($0.966$), while reducing inference time to $7.1$s, corresponding to a $24\times$ speedup. On Huh7, CAP matches the best reported $\mathrm{TRA}$ ($0.966$) with the shortest inference time of $3.2$s.
As for the extreme instability of the EmbedTrack and EBB methods on Huh7 with a $\mathrm{TRA}$ of $0.187$ and $0.626$, respectively, our analysis suggests that the observed results stem from the method’s excessive reliance on the segmentation (SEG) stage and the usage of refined masks during the training of the segmentation model, while Huh7 contains only SEG golden truth (GT) masks without the more refined SEG silver truth (ST). This further substantiates an end-to-end one-stage cell tracking framework, which bypasses the segmentation or detection stage, enhances model stability, and raises the overall performance limits achievable by the model.
Figure 6: Result of Tracking Accuracy and Inference Efficiency across Multiple Benchmark Datasets. CAP consistently achieves the best trade-off between accuracy and efficiency, delivering competitive or superior performance while reducing inference time.
Overall, as shown in Fig. 6, the CAP method demonstrates promising performance and constant high efficiency on inference and label usage across all datasets, consistently achieving high $\mathrm{TRA}$ scores and the lowest $\mathrm{IT}$s. Thus, CAP is a highly effective, efficient, and reliable method for cell tracking in a practical biomedical field.
Figure 8: Cell Lineage (Divisions). Utilizing the proposed Cell Point Trajectory and Visibility representation illustrated in Fig. 2, CAP is able to correctly identify the cell division and track all cells in the sequence. M denotes the mother cell, while D1 and D2 represent daughter cell #1 and daughter cell #2.
4.3 Qualitative Result
To further validate the performance of CAP, we provide qualitative results illustrating two core aspects of cell tracking: cell trajectory estimation and lineage (division) reconstruction.
AEG Sampling U373 GOWT1 Huh7 HeLa PC-3 3T3 RAW264
@@ 4.3 Qualitative Result
0.556 0.941 0.942 0.898 0.908 0.883 0.921 0.878 0.952 0.924 0.829 0.804 0.828 0.843
0.985 0.960 0.966 0.962 0.958 0.921 0.900 0.891 0.980 0.924 0.832 0.824 0.868 0.900
Fig. 7 visualizes representative examples of cell trajectories across multiple time points. Each cell is assigned a distinct color, enabling clear visualization of its movement path. The results demonstrate CAP’s ability to robustly maintain cell identities over extended sequences, even in the presence of densely packed regions, complex motion patterns, and various motion speeds. This consistency highlights CAP’s strength in long-term trajectory preservation and cross-frame association accuracy.
Fig. 8 illustrates the model’s performance on capturing cell division events. CAP successfully identifies division points and accurately tracks the transition from mother cells (M) to daughter cells (D1 and D2), ensuring continuity across frames. The highlighted examples show that CAP not only maintains the mother–daughter relationships but also effectively disambiguates multiple simultaneous divisions within crowded environments. These results underscore the model’s capacity to reconstruct biologically meaningful lineage information, which is critical for downstream analysis in developmental biology and cancer research.
4.4 Ablation Study
We performed extensive experiments to assess the effectiveness of each key component in the CAP framework on both the DeepCell and ISBI CTC datasets: the AEG sampling strategy, the cell joint tracking structure, and the feature stride of the model. To more clearly observe the influence of each key component on individual test sequences, we report the $\mathrm{TRA}$ results for each sequence in the DeepCell datasets.
Effectiveness of AEG Sampling.
To address the imbalance of cell division events during training, we propose the Adaptive Event-Guided (AEG) sampling strategy (Section 3.2.1), which calibrates the training process by ensuring balanced exposure to complete division sequences. As shown in Table 4, disabling AEG sampling leads to substantial performance degradation across datasets. With AEG sampling enabled, the model achieves consistent improvements: for instance, on the HeLa dataset, the best $\mathrm{TRA}$ score rises from $0.898$ to $0.962$, while on U373 it improves from $0.556$ to $0.985$. Similar gains are observed on PC-3 and RAW264. These results confirm that AEG sampling effectively balances event occurrences and significantly enhances the model’s ability to learn from imbalanced data.
Attention U373 GOWT1 Huh7 HeLa PC-3 3T3 RAW264
@@ 4.4 Ablation Study
Effectiveness of Cell Joint Tracking.
We evaluate the role of joint tracking by replacing cross-trajectory attention layers with time-only attention while maintaining model size. As reported in Table 5, this modification consistently reduces performance, with $\mathrm{TRA}$ drops ranging from $0.015$ to $0.055$ across datasets. Incorporating both time and joint attention yields notable improvements, particularly on HeLa where the $\mathrm{TRA}$ increases from $0.853$ to $0.962$ in the best case. This demonstrates that the dual-attention design not only strengthens temporal coherence but also captures relational dependencies between trajectories, leading to more robust tracking in complex cellular environments.
Stride U373 GOWT1 Huh7 HeLa PC-3 3T3 RAW264
@@
Ablation on Feature Stride.
Table 6 highlights the effect of feature stride on model accuracy. Reducing the stride from $8$ to $4$ substantially improves performance by providing higher-resolution features. For example, on the HeLa dataset, the best $\mathrm{TRA}$ score increases from $0.884$ to $0.962$, while on U373 the score rises from $0.531$ to $0.985$. Similar trends are observed across 3T3 and RAW264, demonstrating that higher-resolution features are critical for capturing fine-grained cell dynamics and improving tracking accuracy, especially in challenging datasets with small or densely packed cells. Consequently, we adopt $s=4$ as the default stride setting in CAP framework.
5 Conclusion and Future Work
In this paper, we introduce the one-stage CAP (Cell as Point) framework, which eliminates the need for a separate segmentation or detection stage, which is a significant innovation in cell tracking. This approach substantially improves efficiency by reducing inference time by a factor of $8$ to $32$ and minimizing the need for high-quality data annotation. CAP effectively captures complex cell states and tracks cells in one stage while maintaining high efficiency. Additionally, it addresses common challenges in cell sequence data, such as cell event imbalance and long sequence inference. In conclusion, CAP not only offers considerable practical value in real-world applications due to its efficiency but also pioneers a novel approach in the field of cell tracking.
Limitations and Future Work.
The limitation of our study arises from the nature of the datasets used for evaluation. These datasets were annotated by different experts, and variations in annotation practices may lead to differences in the location of the cell points that CAP generated. Such inconsistencies can introduce discrepancies in performance across datasets, as the model could be sensitive to variations in label quality and annotation style. In addition, differences in imaging devices and cell densities across datasets may further amplify these effects, contributing to dataset-specific performance variation. Looking forward, a promising direction is to design strategies that mitigate the impact of heterogeneous annotations and dataset-specific biases, thereby reducing sensitivity to labeling differences. Consequently, extending CAP toward a more versatile universal model that can generalize effectively across diverse datasets and imaging conditions deserves future exploration.